# BindWeave: Subject-Consistent Video Generation via Cross-Modal Integration

### v1

@@ 1 Introduction
Recent advances in diffusion models Ho et al. (2020); Peebles & Xie (2023); Wang et al. (2025) have enabled significant breakthroughs in video generation Wan et al. (2025); Yang et al. (2024); Hu et al. (2025); Kong et al. ; Polyak et al. (2024); Zheng et al. (2024), achieving outstanding performance on various tasks ranging from text-to-video (T2V) HaCohen et al. (2024); Wan et al. (2025); Chen et al. (2025a) to image-to-video (I2V) Blattmann et al. (2023); Mao et al. (2025). Foundation models such as Wan Wan et al. (2025) and HunyuanVideo Kong et al. now demonstrate the ability to produce high-fidelity, long-duration, and content-rich videos, showcasing immense technological potential. However, despite these advances in visual quality, their practical utility remains constrained by limited controllability. Specifically, current models struggle to exert precise and stable control over key elements within a video, such as the identity of a specific person, the appearance of an object, or a brand logo. This lack of controllability constitutes a core limitation that significantly impedes deployment in customized applications, including personalized content creation, brand marketing, pre-visualization, and virtual try-on.
To address the above challenges, subject-to-video (S2V) Liu et al. (2025) has garnered increasing attention. The core objective of S2V is to ensure that one or more subjects within a video maintain high fidelity in their identity and appearance with respect to the given reference images throughout the entire dynamic sequence. This capability directly addresses the controllability shortcomings of existing general-purpose models, making it possible to generate customized videos based on user-provided subjects. To achieve subject-consistent video generation, some existing works Yuan et al. (2024b); Chen et al. (2025b); Huang et al. (2025); Liu et al. (2025); Jiang et al. (2025) extend a video foundation model to accept multiple reference images as conditioning input. For instance, Phantom Liu et al. (2025) introduces a dual-branch architecture to separately process the text prompt and reference images, subsequently injecting the resulting features into the attention layers of a Diffusion Transformer (DiT) Peebles & Xie (2023) as conditioning. VACE Jiang et al. (2025) designs a video condition unit to unify inputs (text, image/video references, mask) into a unified format, then inject these context signals via residual blocks to guide video generation.
@@
To overcome this bottleneck, we propose BindWeave, a novel framework designed for subject-consistent video generation. To establish the deep cross-modal semantic associations, BindWeave leverages a Multimodal Large Language Model (MLLM) as an intelligent instruction parser to replace the conventional shallow fusion mechanism. Specifically, we first construct a unified, interleaved sequence from reference images and text prompt. This sequence is then processed by a pre-trained MLLM to parse complex spatio-temporal relationships and bind textual commands to their corresponding visual entities. Through this process, the MLLM generates a set of hidden states encoding both the precise identity of each subject and their prescribed interactions. These hidden states then serve as conditioning inputs to our generator, bridging high-level parsing with diffusion-based generation. To provide subject-grounded semantic anchors and further reduce identity drift, we also incorporate CLIP Radford et al. (2021b) features from the reference images. Accordingly, our DiT Peebles & Xie (2023) based generator is jointly conditioned on these hidden states and CLIP features. Together, these conditioning inputs provide comprehensive relational and semantic guidance. To preserve fine-grained appearance details, we augment the video latents during diffusion with VAE Esser et al. (2021) features extracted from the reference images. This collective conditioning on high-level reasoning, semantic identity, and low-level detail ensures the generation of videos with exceptional fidelity and consistency.
We conduct a comprehensive evaluation of BindWeave on the fine-grained opens2v Yuan et al. (2025) benchmark against a diverse set of existing approaches, including leading open-source methods and commercial models. The evaluation assesses key aspects such as subject consistency, temporal naturalness, and text-video alignment. Extensive experiments demonstrate that BindWeave achieves state-of-the-art performance, consistently outperforming all competing methods in subject-consistent video generation. Qualitative results, illustrated in Figure 1, further demonstrate the superior quality of the generated samples. These findings highlight BindWeave’s effectiveness in subject-consistent video generation and its potential as a high-performing solution for both research and commercial applications.
@@ 2.1 Video generation model
Diffusion models have enabled remarkable advancements in video generation. Early methods Singer et al. (2022); Blattmann et al. (2023); Guo et al. (2023) often extended text-to-image models Rombach et al. (2022) for video generation by incorporating temporal modeling modules. More recently, the Diffusion Transformer (DiT) Peebles & Xie (2023) architecture, motivated by its excellent scaling properties, has inspired a new wave of models like Wan Wan et al. (2025), HunyuanVideo Kong et al. (2024a), and Goku Chen et al. (2025a). However, these models focus on general-purpose video generation, and there is still considerable room for improvement in achieving fine-grained control.
@@ 2.2 Subject-consistent video generation
To achieve more fine-grained control, subject-consistent video generation has garnered significant attention. Initial approaches often rely on per-subject optimization, where a pre-trained model is fine-tuned on images of a specific subject, as seen in methods like CustomVideo Wang et al. (2024) and DisenStudio Chen et al. (2024). While effective, this instance-specific tuning is computationally expensive and poses challenges for real-time applications. More recent works have shifted towards end-to-end methods that use conditioning networks or adapters to inject identity information during inference, allowing for generalization to new subjects without retraining. These models, such as IDAnimator He et al. (2024) and ConsisID Yuan et al. (2024a), initially focused on preserving facial identity. This capability was later extended to arbitrary objects and multiple subjects by works like ConceptMaster Huang et al. (2025), SkyReels-A2 Fei et al. (2025),Phantom Liu et al. (2025), and VACE Jiang et al. (2025). Despite this progress, significant challenges remain, particularly maintaining distinct identities and modeling complex interactions in multi-subject scenes.
@@ 3.1 Preliminaries
Diffusion Transformer Models for Text-to-Video Generation. Transformer-based text-to-video diffusion models have shown substantial promise for video content generation. Our BindWeave builds upon a Transformer-based latent diffusion architecture that employs a spatio-temporal Variational Autoencoder (VAE) Wan et al. (2025) to map videos from the pixel level to a compact latent space, where the generative process is performed. Each Transformer block comprises spatiotemporal self-attention, text cross-attention, a time-conditioning MLP, and a feed-forward network (FFN). The cross-attention is conditioned on a text prompt embedding $c_{\text{text}}$ obtained from a T5 encoder $\mathcal{E}_{\text{T5}}$ Raffel et al. (2020). We employ Rectified Flow Liu et al. (2022); Esser et al. (2024) to define the diffusion dynamics, which enables stable training via ordinary differential equations (ODEs) while maintaining equivalence to maximum likelihood objectives. In the forward process of training, random noise is add to clean data $z_{0}$ to generate $z_{t}=(1-t)z_{0}+t\epsilon$, where $\epsilon$ is sampled from a standard normal distribution, $\mathcal{N}(0,I)$, and the time coefficient $t$ is sampled from $[0,1]$. Accordingly, the learning objective becomes the estimation of ground truth velocity field $v_{t}=dz_{t}/dt=\epsilon-z_{0}$. The network $u_{\Theta}$ is trained to this end using the Flow Matching loss Esser et al. (2024):
@@
Natural language offers an accessible interface for diffusion-based video synthesis, yet it often under-specifies subject identity and spatial layout. This motivates the incorporation of a reference image to anchor appearance and geometry in text-to-video pipelines Wan et al. (2025); Kong et al. (2024b). For instance, Wan Wan et al. (2025)injects image information in two ways: first at the input level for fine-grained spatial detail, and second within the cross-attention mechanism for semantic guidance. First, to preserve fine-grained appearance, the reference image $\mathcal{I}_{img}$ is encoded by a VAE into a spatial latent $z_{img}$. This latent is concatenated with the current noisy video latent $\mathbf{x}_{t}$ along the channel dimension. The combined latent is then patchified and linearly embedded to form the initial sequence of tokens for the DiT block:
@@ 3.3 Intelligent Instruction Planning via MLLM
To effectively foster joint cross-modal learning between the text prompt and reference images, we introduces a unified multimodal parsing strategy. Given a text prompt $\mathcal{T}$ and $K$ user-specified subjects, each with a reference image $I_{k}$, we constructs a multimodal sequence $\mathcal{X}$ by appending one image placeholder for each reference image after the text prompt. The MLLM is then provided with this sequence along with the corresponding list of images $\mathcal{I}$:
@@
While this MLLM-derived condition provides invaluable high-level, cross-modal reasoning, we recognize that diffusion models are also highly optimized to interpret fine-grained textual semantics. To provide this complementary signal, we encode the original prompt independently using the T5 text encoder ($\mathcal{E}_{\text{T5}}$)  Raffel et al. (2020) to produce a dedicated textual embedding $c_{\text{text}}$:
@@ 3.5 Training and Inference
Our training data is curated from the 5 million publicly available OpenS2V-5M dataset Yuan et al. (2025). Through a series of filtering strategies, we distill a final, high-quality dataset of approximately 1 million video-text pairs. The training process then follows a two-stage curriculum learning strategy on this data. All training processes are conducted on 512 xPUs, with a global batch size of 512, utilizing a constant learning rate of 5e-6 and the AdamW optimizer. The initial stabilization phase, lasting for approximately 1,000 iterations, utilizes a smaller, core subset selected from the 1 million data for its exceptional quality and representativeness. This initial phase is crucial for adapting the model to the specific demands of the Subject-to-Video (S2V) task, primarily focusing on learning to faithfully preserve a subject’s visual identity while aligning it with the corresponding textual motion commands. This establishes a robust foundation for the subsequent large-scale training. Subsequently, the training transitions to a full-scale phase for an additional 5,000 iterations, where the model is exposed to the entirety of the 1 million curated dataset. This second stage allows the model to build upon its stable foundation and learn from a broader range of high-quality examples, significantly enhancing its generative capabilities and generalization performance.
Inference settings. During inference, our BindWeave accepts a flexible number of reference images (typically 1-4), while a text prompt steers the generation by describing the desired scene and behaviors. Similar to Phantom Liu et al. (2025), we use a prompt rephraser during inference to ensure the text accurately describes the provided reference images. Generation is performed over 50 steps using a rectified flow Liu et al. (2022) trajectory, guided by Classifier-Free Guidance (CFG) Ho & Salimans (2022) with a scale of $\omega$. The guided noise estimate at each step $t$ is computed as:
@@ 4.1 Experimental Settings
Benchmark and Evaluation Metrics. To ensure a fair comparison, we adopt the OpenS2V-Eval benchmark Yuan et al. (2025) and follow its official evaluation protocol, which provides fine-grained assessments of subject consistency and identity fidelity for subject-to-video generation. The benchmark comprises 180 prompts in seven distinct categories, covering scenarios from single-subject (face, body, entity) to multi-subject and human–entity interactions. To quantify performance, we report the protocol’s automated metrics, with higher scores indicating better results across all metrics. These include Aesthetics christophschuhmann (2024) for visual appeal, MotionSmoothness Bradski et al. (2000) for temporal smoothness, MotionAmplitude Bradski et al. (2000) for motion magnitude, and FaceSim Yuan et al. (2024a) for identity preservation. We also use three metrics introduced by OpenS2V-Eval Yuan et al. (2025) that correlate highly with human perception: NexusScore (subject consistency), NaturalScore (naturalness), and GmeScore (text–video relevance).
Implementation details. BindWeave is fine-tuned from a foundation video generation model based on DiT architecture Wan et al. (2025). The T2V and I2V pre-training stages are excluded from this evaluation. For the core instruction planning module, we employ Qwen2.5-VL-7B Bai et al. (2025) as our Multimodal Large Language Model (MLLM). To align the multimodal control signal with the DiT conditioning space, we introduce a lightweight connector that projects the Qwen2.5-VL hidden states. Specifically, the connector features a two-layer MLP with GELU activation. We train our model using the Adam optimizer with a 5e-6 learning rate and a global batch size of 512. To mitigate copy-paste artifacts, we apply data augmentations (e.g., random rotation, scaling) to reference images. During inference, we use 50 denoising steps set the CFG guidance scale $\omega$ to 5.
Baselines. We compare BindWeave with the state-of-the-art video customization methods, including open-sourced methods (Phantom Liu et al. (2025), VACE Jiang et al. (2025), SkyReels-A2 Fei et al. (2025), MAGREF Deng et al. (2025)) and commercial products (Kling-1.6 Kwai (2024), Vidu-2.0 Bao et al. (2024), Pika Lab (2024), Hailuo Team (2024)).
@@ 4.2 Quantitative results
We conduct a comprehensive comparison on the OpenS2V-Eval benchmark Yuan et al. (2025), as shown in Table 1, providing a broad and rigorous evaluation across diverse scenarios. Following the benchmark’s protocol, each method generates 180 videos for evaluation to ensure statistical reliability and coverage of all categories. We report eight automatic metrics as described in Section 4.1 to ensure comprehensive assessment, thereby capturing visual quality, temporal behavior, and semantic alignment in a unified manner. As shown in Table 1, our BinWeave achieves a new state of the art on the overall Total Score, with notably stronger NexusScore that highlights its advantage on subject consistency. Notably, NexusScore Yuan et al. (2025) is designed to address the limitations of prior global-frame CLIP Radford et al. (2021a) or DINO Oquab et al. (2023) comparisons and provide a semantically grounded, noise-resilient assessment that better reflects perceptual identity fidelity. It achieves this via a detect-then-compare strategy that first localizes the true target, crops the relevant regions to suppress background interference, and then computes similarity within a retrieval-based multimodal feature space with confidence and text–image gating, finally aggregating scores over the verified crops for a reliable summary. Importantly, BinWeave also maintains strong competitiveness on other metrics, including FaceSim, Aesthetics, GmeScore, motion-related measures such as MotionSmoothness and MotionAmplitude, and NaturalScore, which respectively reflect its strengths in identity preservation, visual appeal, text–video alignment, temporal coherence and motion magnitude, and overall naturalness across a wide range of prompts and categories.

### v2

@@ 1 Introduction
Recent advances in diffusion models (Ho et al., 2020; Peebles and Xie, 2023; Wang et al., 2025) have enabled significant breakthroughs in video generation (Wan et al., 2025; Yang et al., 2024; Hu et al., 2025; Kong et al., ; Polyak et al., 2024; Zheng et al., 2024; Wu et al., 2025), achieving outstanding performance on various tasks ranging from text-to-video (T2V) (HaCohen et al., 2024; Wan et al., 2025; Chen et al., 2025a) to image-to-video (I2V) (Blattmann et al., 2023; Mao et al., 2025). Foundation models such as Wan (Wan et al., 2025) and HunyuanVideo (Kong et al., ) now demonstrate the ability to produce high-fidelity, long-duration, and content-rich videos, showcasing immense technological potential. However, despite these advances in visual quality, their practical utility remains constrained by limited controllability. Specifically, current models struggle to exert precise and stable control over key elements within a video, such as the identity of a specific person, the appearance of an object, or a brand logo. This lack of controllability constitutes a core limitation that significantly impedes deployment in customized applications, including personalized content creation, brand marketing, pre-visualization, and virtual try-on.
To address the above challenges, subject-to-video (S2V) (Liu et al., 2025) has garnered increasing attention. The core objective of S2V is to ensure that one or more subjects within a video maintain high fidelity in their identity and appearance with respect to the given reference images throughout the entire dynamic sequence. This capability directly addresses the controllability shortcomings of existing general-purpose models, making it possible to generate customized videos based on user-provided subjects. To achieve subject-consistent video generation, some existing works (Yuan et al., 2024b; Chen et al., 2025b; Huang et al., 2025; Liu et al., 2025; Jiang et al., 2025) extend a video foundation model to accept multiple reference images as conditioning input. For instance, Phantom (Liu et al., 2025) introduces a dual-branch architecture to separately process the text prompt and reference images, subsequently injecting the resulting features into the attention layers of a Diffusion Transformer (DiT) (Peebles and Xie, 2023) as conditioning. VACE (Jiang et al., 2025) designs a video condition unit to unify inputs (text, image/video references, mask) into a unified format, then inject these context signals via residual blocks to guide video generation.
@@
To overcome this bottleneck, we propose BindWeave, a novel framework designed for subject-consistent video generation. To establish the deep cross-modal semantic associations, BindWeave leverages a Multimodal Large Language Model (MLLM) as an intelligent instruction parser to replace the conventional shallow fusion mechanism. Specifically, we first construct a unified, interleaved sequence from reference images and text prompt. This sequence is then processed by a pre-trained MLLM to parse complex spatio-temporal relationships and bind textual commands to their corresponding visual entities. Through this process, the MLLM generates a set of hidden states encoding both the precise identity of each subject and their prescribed interactions. These hidden states then serve as conditioning inputs to our generator, bridging high-level parsing with diffusion-based generation. To provide subject-grounded semantic anchors and further reduce identity drift, we also incorporate CLIP (Radford et al., 2021b) features from the reference images. Accordingly, our DiT (Peebles and Xie, 2023) based generator is jointly conditioned on these hidden states and CLIP features. Together, these conditioning inputs provide comprehensive relational and semantic guidance. To preserve fine-grained appearance details, we augment the video latents during diffusion with VAE (Esser et al., 2021) features extracted from the reference images. This collective conditioning on high-level reasoning, semantic identity, and low-level detail ensures the generation of videos with exceptional fidelity and consistency.
We conduct a comprehensive evaluation of BindWeave on the fine-grained OpenS2V (Yuan et al., 2025) benchmark against a diverse set of existing approaches, including leading open-source methods and commercial models. The evaluation assesses key aspects such as subject consistency, temporal naturalness, and text-video alignment. Extensive experiments demonstrate that BindWeave achieves state-of-the-art performance, consistently outperforming all competing methods in subject-consistent video generation. Qualitative results, illustrated in Figure 1, further demonstrate the superior quality of the generated samples. These findings highlight BindWeave’s effectiveness in subject-consistent video generation and its potential as a high-performing solution for both research and commercial applications.
@@ 2.1 Video generation model
Diffusion models have enabled remarkable advancements in video generation. Early methods (Singer et al., 2022; Blattmann et al., 2023; Guo et al., 2023) often extended text-to-image models (Rombach et al., 2022) for video generation by incorporating temporal modeling modules. More recently, the Diffusion Transformer (DiT) (Peebles and Xie, 2023) architecture, motivated by its excellent scaling properties, has inspired a new wave of models like Wan (Wan et al., 2025), HunyuanVideo (Kong et al., 2024a), and Goku (Chen et al., 2025a). However, these models focus on general-purpose video generation, and there is still considerable room for improvement in achieving fine-grained control.
@@ 2.2 Subject-consistent video generation
To achieve more fine-grained control, subject-consistent video generation has garnered significant attention. Initial approaches often rely on per-subject optimization, where a pre-trained model is fine-tuned on images of a specific subject, as seen in methods like CustomVideo (Wang et al., 2024) and DisenStudio (Chen et al., 2024). While effective, this instance-specific tuning is computationally expensive and poses challenges for real-time applications. More recent works have shifted towards end-to-end methods that use conditioning networks or adapters to inject identity information during inference, allowing for generalization to new subjects without retraining. These models, such as IDAnimator (He et al., 2024) and ConsisID (Yuan et al., 2024a), initially focused on preserving facial identity. This capability was later extended to arbitrary objects and multiple subjects by works like ConceptMaster (Huang et al., 2025), SkyReels-A2 (Fei et al., 2025),Phantom (Liu et al., 2025), and VACE (Jiang et al., 2025). Despite this progress, significant challenges remain, particularly maintaining distinct identities and modeling complex interactions in multi-subject scenes.
@@ 3.1 Preliminaries
Diffusion Transformer Models for Text-to-Video Generation. Transformer-based text-to-video diffusion models have shown substantial promise for video content generation. Our BindWeave builds upon a Transformer-based latent diffusion architecture that employs a spatio-temporal Variational Autoencoder (VAE) (Wan et al., 2025) to map videos from the pixel level to a compact latent space, where the generative process is performed. Each Transformer block comprises spatiotemporal self-attention, text cross-attention, a time-conditioning MLP, and a feed-forward network (FFN). The cross-attention is conditioned on a text prompt embedding $c_{\text{text}}$ obtained from a T5 encoder $\mathcal{E}_{\text{T5}}$ (Raffel et al., 2020). We employ Rectified Flow (Liu et al., 2022; Esser et al., 2024) to define the diffusion dynamics, which enables stable training via ordinary differential equations (ODEs) while maintaining equivalence to maximum likelihood objectives. In the forward process of training, random noise is add to clean data $z_{0}$ to generate $z_{t}=(1-t)z_{0}+t\epsilon$, where $\epsilon$ is sampled from a standard normal distribution, $\mathcal{N}(0,I)$, and the time coefficient $t$ is sampled from $[0,1]$. Accordingly, the learning objective becomes the estimation of ground truth velocity field $v_{t}=dz_{t}/dt=\epsilon-z_{0}$. The network $u_{\Theta}$ is trained to this end using the Flow Matching loss (Esser et al., 2024):
@@
Natural language offers an accessible interface for diffusion-based video synthesis, yet it often under-specifies subject identity and spatial layout. This motivates the incorporation of a reference image to anchor appearance and geometry in text-to-video pipelines Wan et al. (2025); Kong et al. (2024b). For instance, Wan (Wan et al., 2025)injects image information in two ways: first at the input level for fine-grained spatial detail, and second within the cross-attention mechanism for semantic guidance. First, to preserve fine-grained appearance, the reference image $\mathcal{I}_{img}$ is encoded by a VAE into a spatial latent $z_{img}$. This latent is concatenated with the current noisy video latent $\mathbf{x}_{t}$ along the channel dimension. The combined latent is then patchified and linearly embedded to form the initial sequence of tokens for the DiT block:
@@ 3.3 Intelligent Instruction Planning via MLLM
To effectively foster joint cross-modal learning between the text prompt and reference images, we introduce a unified multimodal parsing strategy. Given a text prompt $\mathcal{T}$ and $K$ user-specified subjects, each with a reference image $I_{k}$, we constructs a multimodal sequence $\mathcal{X}$ by appending one image placeholder for each reference image after the text prompt. The MLLM is then provided with this sequence along with the corresponding list of images $\mathcal{I}$:
@@
While this MLLM-derived condition provides invaluable high-level, cross-modal reasoning, we recognize that diffusion models are also highly optimized to interpret fine-grained textual semantics. To provide this complementary signal, we encode the original prompt independently using the T5 text encoder ($\mathcal{E}_{\text{T5}}$)  (Raffel et al., 2020) to produce a dedicated textual embedding $c_{\text{text}}$:
@@ 3.5 Training and Inference
Our training data is curated from the 5 million publicly available OpenS2V-5M dataset (Yuan et al., 2025). Through a series of filtering strategies, we distill a final, high-quality dataset of approximately 1 million video-text pairs. The training process then follows a two-stage curriculum learning strategy on this data. All training processes are conducted on 512 xPUs, with a global batch size of 512, utilizing a constant learning rate of 5e-6 and the AdamW optimizer. The initial stabilization phase, lasting for approximately 1,000 iterations, utilizes a smaller, core subset selected from the 1 million data for its exceptional quality and representativeness. This initial phase is crucial for adapting the model to the specific demands of the Subject-to-Video (S2V) task, primarily focusing on learning to faithfully preserve a subject’s visual identity while aligning it with the corresponding textual motion commands. This establishes a robust foundation for the subsequent large-scale training. Subsequently, the training transitions to a full-scale phase for an additional 5,000 iterations, where the model is exposed to the entirety of the 1 million curated dataset. This second stage allows the model to build upon its stable foundation and learn from a broader range of high-quality examples, significantly enhancing its generative capabilities and generalization performance.
Inference settings. During inference, our BindWeave accepts a flexible number of reference images (typically 1-4), while a text prompt steers the generation by describing the desired scene and behaviors. Similar to Phantom (Liu et al., 2025), we use a prompt rephraser during inference to ensure the text accurately describes the provided reference images. Generation is performed over 50 steps using a rectified flow Liu et al. (2022) trajectory, guided by Classifier-Free Guidance (CFG) Ho and Salimans (2022) with a scale of $\omega$. The guided noise estimate at each step $t$ is computed as:
@@ 4.1 Experimental Settings
Benchmark and Evaluation Metrics. To ensure a fair comparison, we adopt the OpenS2V-Eval benchmark (Yuan et al., 2025) and follow its official evaluation protocol, which provides fine-grained assessments of subject consistency and identity fidelity for subject-to-video generation. The benchmark comprises 180 prompts in seven distinct categories, covering scenarios from single-subject (face, body, entity) to multi-subject and human–entity interactions. To quantify performance, we report the protocol’s automated metrics, with higher scores indicating better results across all metrics. These include Aesthetics (christophschuhmann, 2024) for visual appeal, MotionSmoothness (Bradski et al., 2000) for temporal smoothness, MotionAmplitude (Bradski et al., 2000) for motion magnitude, and FaceSim (Yuan et al., 2024a) for identity preservation. We also use three metrics introduced by OpenS2V-Eval (Yuan et al., 2025) that correlate highly with human perception: NexusScore (subject consistency), NaturalScore (naturalness), and GmeScore (text–video relevance).
Implementation details. BindWeave is fine-tuned from a foundation video generation model based on DiT architecture (Wan et al., 2025). The T2V and I2V pre-training stages are excluded from this evaluation. For the core instruction planning module, we employ Qwen2.5-VL-7B (Bai et al., 2025) as our Multimodal Large Language Model (MLLM). To align the multimodal control signal with the DiT conditioning space, we introduce a lightweight connector that projects the Qwen2.5-VL hidden states. Specifically, the connector features a two-layer MLP with GELU activation. We train our model using the Adam optimizer with a 5e-6 learning rate and a global batch size of 512. To mitigate copy-paste artifacts, we apply data augmentations (e.g., random rotation, scaling) to reference images. During inference, we use 50 denoising steps set the CFG guidance scale $\omega$ to 5.
Baselines. We compare BindWeave with the state-of-the-art video customization methods, including open-sourced methods (Phantom (Liu et al., 2025), VACE (Jiang et al., 2025), SkyReels-A2 (Fei et al., 2025), MAGREF (Deng et al., 2025)) and commercial products (Kling-1.6 (Kwai, 2024), Vidu-2.0 (Bao et al., 2024), Pika (Lab, 2024), Hailuo (Team, 2024)).
@@ 4.2 Quantitative results
We conduct a comprehensive comparison on the OpenS2V-Eval benchmark (Yuan et al., 2025), as shown in Table 1, providing a broad and rigorous evaluation across diverse scenarios. Following the benchmark’s protocol, each method generates 180 videos for evaluation to ensure statistical reliability and coverage of all categories. We report eight automatic metrics as described in Section 4.1 to ensure comprehensive assessment, thereby capturing visual quality, temporal behavior, and semantic alignment in a unified manner. As shown in Table 1, our BindWeave achieves a new state of the art on the overall Total Score, with notably stronger NexusScore that highlights its advantage on subject consistency. Notably, NexusScore (Yuan et al., 2025) is designed to address the limitations of prior global-frame CLIP (Radford et al., 2021a) or DINO (Oquab et al., 2023) comparisons and provide a semantically grounded, noise-resilient assessment that better reflects perceptual identity fidelity. It achieves this via a detect-then-compare strategy that first localizes the true target, crops the relevant regions to suppress background interference, and then computes similarity within a retrieval-based multimodal feature space with confidence and text–image gating, finally aggregating scores over the verified crops for a reliable summary. Importantly, BindWeave also maintains strong competitiveness on other metrics, including FaceSim, Aesthetics, GmeScore, motion-related measures such as MotionSmoothness and MotionAmplitude, and NaturalScore, which respectively reflect its strengths in identity preservation, visual appeal, text–video alignment, temporal coherence and motion magnitude, and overall naturalness across a wide range of prompts and categories.